# Qualitative Study for LLM-assisted Design Study Process: Strategies, Challenges, and Roles

### v2

@@ Abstract
Design studies aim to create visualization solutions for real-world problems of different application domains. Recently, the emergence of large language models (LLMs) has introduced new opportunities to enhance the design study process, providing capabilities such as creative problem-solving, data handling, and insightful analysis. However, despite their growing popularity, there remains a lack of systematic understanding of how LLMs can effectively assist researchers in visualization-specific design studies. In this paper, we conducted a multi-stage qualitative study to fill this gap, involving 30 design study researchers from diverse backgrounds and expertise levels. Through in-depth interviews and carefully-designed questionnaires, we investigated strategies for utilizing LLMs, the challenges encountered, and the practices used to overcome them. We further compiled and summarized the roles that LLMs can play across different stages of the design study process. Our findings highlight practical implications to inform visualization practitioners, and provide a framework for leveraging LLMs to enhance the design study process in visualization research.
@@
An overview of the summarized roles of LLMs in the LLM-assisted design study process. Following the framework proposed by Sedlmair et al. [34], we coded all nine stages of the design study by the compiled roles: 1) Connector: LLMs are used to bridge gaps between visualization researchers and domain experts; 2) Assistant: LLMs are commonly used for repetitive tasks to enhance productivity; 3) Simulator: LLMs support simulating user behavior, predicting feedback, and testing workflows to uncover usability issues and hidden use cases; 4) Programmer: LLMs are asked to assist in coding tasks, from prototyping to debugging. The bottom shows the 30 researchers’ ratings, of importance, difficulty, necessity of using LLMs, and difficulty of using LLMs for each stage.
@@ Introduction
Sedlmair et al. introduced the term design study to describe an applied research methodology that focuses on creating visualizations to address specific, real-world problems [34]. More specifically, they defined a visualization design study as “a project in which visualization researchers analyze a specific problem faced by domain experts and design a visualization system that supports solving this problem” [34]. Building upon this definition, a nine-stage methodology framework was proposed for conducting design studies, as shown in Fig. Qualitative Study for LLM-assisted Design Study Process: Strategies, Challenges, and Roles, which has become a common guide for this category of research in the visualization community. Design studies are crucial because they bridge the gap between theoretical visualization research and practical applications, ensuring that visualization tools are both usable and useful in real-world contexts [29]. Basically, researchers conduct design studies through labor-intensive processes (e.g., connecting with domain experts, extracting requirements via interviews, and iteratively designing and developing visualization systems), which requires significant experience and refined knowledge to gain effective solutions.
@@
Meanwhile, Large Language Models (LLMs) have gained increasing popularity in recent years due to their impressive capabilities in natural language processing and generation [48]. LLMs are now widely used to enhance various aspects of the design study process, including assisting in creative problem-solving, handling large volumes of data, and generating insightful analysis results [18, 50, 24, 38]. Thus, by leveraging these capabilities, LLMs show the potential to significantly streamline the workflow of design studies and make the process more efficient and effective [15, 2].
@@
Despite the increasing use of LLMs in conducting design studies, the question of how LLMs can effectively assist researchers in this domain remains underexplored. For example, researchers with limited expertise need to “deepen” their understanding of LLM-assisted design studies to reach higher levels of proficiency. Also, those with a moderate level of expertise can master the practice from peers to quickly “broaden” their knowledge and practice. However, existing research  [36, 12, 14] on how LLMs can aid co-design methodologies in other fields, such as poster design [25], cannot be easily generalized to the field of visualization due to its visualization-specific workflows and methodologies. Thus, a systematic study on how to facilitate the LLM-assisted design study process is urgently needed in the visualization community, which points out several specific challenges for achieving an effective LLM-assisted design study process. First, design study researchers often utilize LLMs based on their own understanding and personal experience, without any systematic and rigorous stage-by-stage guidance as a reference. Second, researchers are often unaware of the potential challenges they may face during the study, which can lead to significant confusion and time-consuming corrective efforts [42]. Even though these challenges have been recognized, individual researchers often lack effective practices to address them. Third, the absence of a high-level summary of LLM characteristics makes it difficult for design study researchers, especially those with less experience, to build a clear mental model of how to effectively use LLMs.
@@
To fill the research gap, we conducted a multi-stage qualitative study that investigated common strategies and challenges for design study researchers using LLMs. To provide a comprehensive perspective, we invited 30 design study developers and researchers from five different countries with varying levels of visualization expertise (i.e., novice, advanced beginner, intermediate, proficient, and expert) to share their insights. To address the first need above, we conducted interviews with those participants to explore their strategies for using LLMs to assist design studies. We then asked them follow-up questions to identify the challenges they encountered while implementing these strategies and also the practices they used to address them. To further characterize the types of tasks that LLMs can actually help with, we performed a post-study analysis and derived a set of roles that LLMs may play during the design study stages by synthesizing all feedback. In addition to the interviews, we also incorporated a questionnaire to gather more nuanced insights into how researchers navigate each stage of the design study methodology [34] via comparing their viewpoints with and without LLM assistance. Finally, we distilled implications from all the findings that can inform both visualization practitioners and LLM researchers of a more productive LLM-assisted design study process in the future.
@@ 1 Related Work
1.1 Human-LLM Collaboration in Visualzation
@@
Despite some insightful knowledge has been yielded from the above studies, they barely investigated the high-level summary of how LLMs are utilized in the whole process of the design study for visualization, which can directly inform design study researchers of the effective usage of LLMs.
@@ 1.3 Qualitative Study in Visualization
To explore the strategies, challenges, and roles of the LLM-assisted design study process, we conducted a multi-stage qualitative study, involving participant profiling, an interview, and a follow-up questionnaire. This section introduced our methodological approach. We first described the participants involved, the data collection methods employed, and the procedures followed throughout the study.
@@ 2.1 Participants
Participants’ expertise in visualization and design study was identified by themselves into five levels: novice, advanced beginner, intermediate, proficient, and expert. Meanwhile, the participants’ design study research spanned a wide range of interdisciplinary domains, including biomedical science, finance, smart cities, sports, etc. In terms of research productivity, our participants had contributed to over 300 research papers in total, of which 179 were accepted in peer-reviewed venues. Novice participants typically had limited publication records, often having contributed to one study. Proficient participants demonstrated more substantial outputs and many of them had authored 10 to 20 papers.
Furthermore, we also requested participants to report the most-used LLM variants and the frequency of LLM usage in their workflows. We gathered this information to understand their overall familiarity with LLMs and the extent to which LLMs are important to their research workflows.
@@ 2.2 Data Collection
We illustrate our data collection process in Fig. 1, as we used the visual representation to better illustrate the study workflow. Firstly, all interviews were conducted via online video calls. Before each interview, we asked if the participant agreed with the recording, and we recorded the whole session upon their consent. At the same time, we took notes and created sketches based on the visualizations described by the participants during each session. We shared these notes and sketches with participants to make sure they accurately reflected participants’ original intentions. Additionally, we documented general impressions and recalled details immediately to capture any points that may not have been noted in real-time.
For the stages of participant profiling (Fig. 1(A)) and questionnaire (Fig. 1(C)), we verbally inquired all interviewees during the meeting right before and after the qualitative interview. Note that all participants consented to provide their demographic information before we moved on to the stage of participant profiling. All data was collected via Google Sheets and every participants can access and edit it anytime.
@@ 2.3 Procedure
Participant profiling. Before the interviews began, participants were asked to provide demographic information, their visualization expertise, and LLM usage frequency. Additionally, participants were asked to indicate their primary design study domain. To further characterize their reliance on LLMs, we asked participants to report their most-used LLM services and the frequency of usage. The LLM usage frequency refers to the approximate number of prompts one sends per day during they conduct the design study process. We categorized the number into four ranges: 0–10 prompts per day, 11–20 prompts per day, 21–30 prompts per day, and more than 30 prompts per day. The participant profiling stage was brief, with each session lasting approximately 5 minutes.
Interview. At the beginning of each interview session, we first introduced the Sedlmair et al. [34]’s nine-stage framework to participants. To ensure clarity, we provided detailed explanations of each stage, supplemented by examples to help participants understand the purpose and process of each stage. After that, we conducted one-on-one, semistructured, hour-long interviews with each participant. The interviews were designed around the three phases outlined in the original paper (as illustrated in Fig. 3), i.e., precondition, core, and analysis. During the interviews, we first asked participants to describe the strategies they typically employ for each phase in their routine tasks. Subsequently, we discussed with participants about the challenges or pitfalls in applying these strategies and how they attempted to address these challenges.
Questionnaire. After completing the interviews, all participants were asked to fill in a post-study questionnaire. This questionnaire included four questions for each individual stage in the nine-stage framework, totaling 36 items (Fig. 3). Specifically, participants were first asked to rate the importance and difficulty of each stage without LLMs. Then they also rated the necessity of each stage requiring the assistance of LLMs, followed by the rating for asking how difficult it is to utilize LLMs in this stage. These ratings were based on a 7-point Likert scale, where 1 indicated “not important/difficult at all” and 7 indicated “extremely important/difficult”. The purpose of the questionnaire was twofold: 1) To understand how much participants value each stage and perceive its difficulty when performing the design study process without LLM assistance; 2) To explore the perceived necessity of using LLMs for each stage, as well as the difficulty of incorporating LLM assistance into the process. We used the responses to this questionnaire to complement the qualitative feedback gathered during the interviews and broaden our understanding of how participants with varying attitudes toward LLMs apply their research strategies in design studies.
@@
Post-study analysis. We conducted a post-interview analysis to identify the high-level roles that LLMs can play in the design study process. The goal was to unify the diverse strategies employed by participants with different backgrounds and experiences. Our analysis mainly focused on the qualitative feedback collected during the interview step. Furthermore, we leveraged the information from participant profiling and their questionnaire responses to more accurately characterize individual researchers according to their expertise, levels of LLM usage, and attitudes toward LLM-assisted workflows. To compile the roles, a team of three researchers independently reviewed the collected qualitative data, including interview transcripts, questionnaire responses, and participant profiles. Each researcher proposed an initial set of roles based on their interpretation of the data. These roles were then discussed in a series of group meetings to reach a consensus. In cases where disagreements arose, the team revisited the original data and engaged in iterative discussions to refine the roles until all members reached an agreement.
@@ 3 Roles of LLMs in Design Studies
In this section, we introduce the roles of LLMs in supporting the design study process. By analyzing the strategies collected from the 30 visualization researchers during the qualitative interviews, we identified four roles: connector, simulator, programmer, and assistant. We first introduce the roles and then present the main strategies related to these roles in Section 4.
@@
Connector: The connector role focuses on bridging the gap between visualization researchers and domain experts. Specifically, LLMs in this role help researchers from the visualization community learn domain-specific knowledge, extract design requirements, and interact with domain experts more effectively. For novice researchers, LLMs serve as a learning tool that provides background knowledge about the application domain, helping them quickly understand target concepts. For proficient PhD students, on the other hand, LLMs assist in analyzing domain-specific workflows, identifying key tasks, and collecting practical design requirements. For example, a novice researcher might use an LLM to understand the basics of bioinformatics before beginning a visualization project (e.g., P6). Meanwhile, a proficient PhD student could use an LLM to record and summarize how bioinformatics researchers analyze tumor tissue slices, ensuring their visualization aligns with real-world tasks. Experts might use LLMs to update their knowledge of a certain domain or generate drafts for academic papers based on high-level summaries. Additionally, LLMs support communication by helping researchers articulate their needs to domain experts and translate domain-specific terminology into accessible language.
@@
Simulator: For the simulator role, LLMs are often used to simulate responses to questionnaires, predict potential user feedback, or test outcomes of workflows. By simulating user behavior, LLMs help researchers uncover hidden use cases, evaluate system usability, and identify cognitive load issues before presenting the system to real users. Additionally, LLMs can take on the role of a critical reviewer, challenging the reasoning and logic behind a system and suggesting potential improvements. For example, a proficient PhD student might use an LLM to simulate user responses during the evaluation process, combining these responses with real user feedback to refine their system. This role is most frequently relied upon by proficient PhD students who are experienced enough to build simulated insights into their design studies effectively.
@@
Programmer: The programmer role is a commonly used applications of LLMs, assisting researchers with a wide range of software development tasks, from prototyping to debugging. LLMs help generate, optimize, and revise code, streamlining the technical implementation process. Specifically, researchers at varying levels of expertise, including novices, intermediates, and proficient researchers, reported using LLMs for programming support. For novices, LLMs provide guidance for setting up their code frameworks or generating simple scripts, such as creating mock datasets or pre-processing data. Intermediate researchers use LLMs for more complex coding tasks, such as generating JavaScript code for implementing visualization designs. Proficient researchers are able to utilize LLMs to improve the overall user interactions of their visualization system prototype. For example, an intermediate researcher could use an LLM to create JavaScript code that translates hand-drawn drafts into a simple visualization. Meanwhile, a proficient researcher might use an LLM to simplify SVG elements or optimize a visualization’s runtime performance.
@@
Assistant: The assistant role refers to the general-purpose support LLMs offer to boost researchers’ productivity by automating repetitive or time-consuming tasks. Common users include searching for academic papers, summarizing a collaborator’s background from a CV, or generating scripts from user study recordings. Unlike the other roles, this role is used by researchers across all experience levels. For example, a novice researcher might use an LLM to summarize a complex academic paper into key points they can easily understand. A postdoc researcher could use it to transcribe interview recordings and distill key takeaways from them.

### v4

@@ Abstract
Design studies aim to develop visualization solutions for real-world problems across various application domains. Recently, the emergence of large language models (LLMs) has introduced new opportunities to enhance the design study process, providing capabilities such as creative problem-solving, data handling, and insightful analysis. However, despite their growing popularity, there remains a lack of systematic understanding of how LLMs can effectively assist researchers in visualization-specific design studies. In this paper, we conducted a multi-stage qualitative study to fill this gap, which involved 30 design study researchers from diverse backgrounds and expertise levels. Through in-depth interviews and carefully-designed questionnaires, we investigated strategies for utilizing LLMs, the challenges encountered, and the practices used to overcome them. We further compiled the roles that LLMs can play across different stages of the design study process. Our findings highlight practical implications to inform visualization practitioners, and also provide a framework for leveraging LLMs to facilitate the design study process in visualization research.
@@
An overview of the summarized roles of LLMs in the LLM-assisted design study process. Following the framework proposed by Sedlmair et al. [34], we coded all nine stages of the design study by the compiled roles: 1) Connector: LLMs are used to bridge gaps between visualization researchers and domain experts; 2) Assistant: LLMs are commonly used for repetitive tasks to enhance productivity; 3) Simulator: LLMs support simulating user behavior, predicting feedback, and testing workflows to uncover usability issues and hidden use cases; 4) Programmer: LLMs are asked to assist in coding tasks, from prototyping to debugging. The bottom section shows ratings from 30 researchers on the importance, difficulty, and necessity of using LLMs, as well as the difficulty when using LLMs at each stage.
@@ Introduction
Sedlmair et al. introduced the term design study to describe an applied research methodology that focuses on creating visualizations to address specific, real-world problems [34]. More specifically, they defined a visualization design study as “a project in which visualization researchers analyze a specific problem faced by domain experts and design a visualization system that supports solving this problem” [34]. Building upon this definition, a nine-stage methodology framework was proposed for conducting design studies, as shown in Fig. Qualitative Study for LLM-assisted Design Study Process: Strategies, Challenges, and Roles, which has become a common guidance for this field of research in the visualization community. Design studies are crucial because they bridge the gap between theoretical visualization research and practical applications, ensuring that visualization tools are both usable and useful in real-world contexts [29]. Basically, researchers conduct design studies through labor-intensive processes (e.g., connecting with domain experts, extracting requirements via interviews, and iteratively designing and developing visualization systems), which require significant effort and refined knowledge to gain effective solutions.
@@
Meanwhile, Large Language Models (LLMs) have shown increasing popularity in recent years due to their impressive problem-solving capabilities with natural language [48]. Thus, LLMs are now widely used to enhance various aspects of the design study process, including assisting in creative problem-solving, handling large volumes of data, and generating insightful findings [19, 50, 25, 38]. By leveraging these advantages, LLMs show the potential to significantly streamline the workflow of design studies and make its process more efficient and effective [16, 2].
@@
Despite the increasing use of LLMs in conducting design studies, the question of how LLMs can effectively assist researchers in this domain remains underexplored. For example, researchers with limited expertise need to “broaden” their practice of LLM-assisted design studies to reach higher levels of proficiency, while those with a moderate or higher level of expertise can grasp the experience from peers to quickly “deepen” their strategies. However, existing research  [36, 13, 15] on how LLMs can aid co-design methodologies in other fields [31, 37, 11, 6, 41, 40, 17] cannot be easily generalized to the field of visualization due to its visualization-specific workflows and methodologies. To this end, a systematic study on how to facilitate the LLM-assisted design study process is urgently needed in the visualization community, which points out several specific challenges for achieving an effective study procedure. First, design study researchers often utilize LLMs in the way on their own understanding and personal experience, without any systematic and rigorous stage-by-stage guidance as a reference. Second, researchers are often unaware of the potential challenges they may face during the study, which can lead to significant confusion and require time-consuming efforts for correction [42]. Even though these challenges have been recognized, individual researchers often lack an effective approach to address them. Third, the absence of a high-level summary of LLM characteristics makes it difficult for design study researchers, especially those with relatively less experience, to build an intuitive mental model of utilizing LLMs effectively.
@@
To fill the research gap, we conducted a multi-stage qualitative study that investigated common strategies and challenges for design study researchers using LLMs. To provide a holistic perspective, we invited 30 design study developers and researchers from five different countries with varying levels of visualization expertise (i.e., novice, advanced beginner, intermediate, proficient, and expert) to share their insights. More specifically, to address the first need above, we first conducted interviews with those participants to explore their strategies for using LLMs for conducting design studies. We then asked them follow-up questions to identify the challenges they encountered while implementing these strategies and also collected the practices they used to address them. To further characterize the types of tasks that LLMs can actually help with, we performed a post-study analysis and derived a set of roles that LLMs may play during the design study stages by synthesizing all the above feedback. In addition to the interviews, we also incorporated a questionnaire to gather more nuanced insights into how researchers navigate each stage of the design study methodology [34] via comparing their viewpoints with and without LLM assistance. Finally, we distilled implications from all the findings that can inform both visualization practitioners and LLM researchers of how to achieve a more productive LLM-assisted design study process in the future.
@@ 1 Related Work
1.1 Human-LLM Collaboration in Visualization
@@
Despite some insightful knowledge being yielded from the above studies, they barely investigated the high-level summary of how LLMs can be used in the process of preparing, performing, and synthesizing design studies in visualization, which can directly inform design study researchers of a more effective usage of LLMs.
@@ 1.3 Qualitative Study in Visualization
To explore the strategies, challenges, and roles of the LLM-assisted design study process, we conducted a multi-stage qualitative study, including participant profiling, an interview, a follow-up questionnaire, and a post-study analysis. This section introduced our methodological approach. We first described the participants involved, the data collection methods employed, and lastly, the procedures followed throughout the study.
@@ 2.1 Participants
Participants’ expertise in visualization and design study was identified by themselves into five levels: novice, advanced beginner, intermediate, proficient, and expert. Meanwhile, the participants’ design study research spanned a wide range of interdisciplinary domains, including biomedical science, finance, smart cities, sports, etc. In terms of research productivity, our participants had contributed to over 300 research papers in total, of which 179 were accepted in peer-reviewed venues. Novice participants typically had limited publication records, often having contributed to one study.
Furthermore, we also requested participants to report the most-used LLM variants and the frequency of LLM usage in their workflows.
@@ 2.2 Data Collection
We illustrate our data collection process in Fig. 1, as we planned to use the visual representation to better illustrate the study workflow. All interviews were conducted via online video calls. Before each interview, we asked if the participant agreed with the recording, and we recorded the whole session upon their consent. At the same time, we took notes and created sketches based on the visualizations described by the participants during each session. We shared these notes and sketches with participants to make sure they accurately reflected participants’ original intentions. Additionally, we documented general impressions and recalled details immediately to capture any points that may not have been noted in real-time.
For the stages of participant profiling (Fig. 1(A)) and questionnaire (Fig. 1(C)), we verbally inquired of all interviewees during the meeting right before and after the qualitative interview. Note that all participants consented to provide their demographic information before we moved on to the stage of participant profiling. All data was collected via Google Sheets and every participants can access and edit it anytime.
@@ 2.3 Procedure
Participant profiling. Before the interviews began, participants were asked to provide demographic information, their visualization expertise and LLM usage frequency. Additionally, participants were asked to indicate their primary design study domain. In addition to the investigation of their demographics, educational background, and visualization study engagement illustrated above, we also asked participants to report their most-used LLM variants and the frequency of usage to characterize their reliance on LLMs. The LLM usage frequency refers to the approximate number of prompts used per day during they conduct the design study process. We categorized the number into four ranges: 0–10 prompts per day, 11–20 prompts per day, 21–30 prompts per day, and more than 30 prompts per day. The participant profiling stage was brief, with each session lasting approximately 5 minutes.
Interview. At the beginning of each interview session, we first introduced the Sedlmair et al. [34]’s nine-stage framework to participants. To ensure clarity, we provided detailed explanations of each stage, supplemented by examples to help participants understand the purpose of each stage. After that, we conducted one-on-one, semistructured, hour-long interviews with each participant. The interviews were designed around the three phases outlined in the original paper (as illustrated in Fig. 3), i.e., precondition, core, and analysis. During the interviews, we first asked participants to describe the strategies they typically employ for each phase in their routine tasks. Subsequently, we discussed with participants about the challenges or pitfalls in applying these strategies and how they attempted to address these challenges.
Questionnaire. After completing the interviews, all participants were asked to fill in a post-study questionnaire. This questionnaire included four questions for each individual stage in the nine-stage framework, totaling 36 items (Fig. 3). Specifically, participants were first asked to rate the importance and difficulty of each stage without LLMs. Then they also rated the necessity of each stage requiring the assistance of LLMs, followed by the rating for asking how difficult it is to utilize LLMs in this stage. These ratings were based on a 7-point Likert scale, where 1 indicated “not important/difficult at all” and 7 indicated “extremely important/difficult”. The purpose of the questionnaire was twofold: 1) To understand how much participants value each stage and perceive its difficulty when performing the design study process without LLM assistance; 2) To explore the perceived necessity of using LLMs for each stage, as well as the difficulty of incorporating LLM assistance into the process. We used the responses to this questionnaire to complement the qualitative feedback gathered during the interviews, making it more insightful for summarizing the strategies and compiling the roles in the subsequent phases.
@@
Post-study analysis. We conducted a post-interview analysis to identify the high-level roles that LLMs can play in the design study process. The goal was to unify the diverse strategies employed by participants with different backgrounds and experiences. We leveraged the information from participant profiling and their questionnaire responses to more accurately characterize individual researchers. To compile the roles, a team of three researchers independently reviewed the collected qualitative data, including interview transcripts, questionnaire responses, and participant profiles. Each researcher proposed an initial set of roles based on their interpretation of the data. These roles were then discussed in a series of group meetings to reach a consensus. In cases where disagreements arose, the team revisited the original data and engaged in iterative discussions to refine the roles until all members reached an agreement.
@@ 3 Roles of LLMs in Design Studies
In this section, we introduce the roles of LLMs in supporting the design study process. By analyzing the strategies collected from the 30 visualization researchers during the qualitative interviews, we identified four roles, i.e., connector, simulator, programmer, and assistant. We first explain the roles briefly and then present the main strategies related to these roles in detail in Section 4.
@@
Connector: The connector role focuses on bridging the gap between visualization researchers and domain experts. More specifically, LLMs in this role help researchers from the visualization community learn domain-specific knowledge, help to extract design requirements, and work with domain experts more effectively. For novice researchers, LLMs basically serve as a learning tool that provides basic background knowledge about the application domain, helping them quickly understand specific concepts. For visualization researchers, on the other hand, LLMs assist in analyzing domain-specific workflows, identifying key tasks, and collecting practical design requirements. For example, a proficient PhD student could use an LLM to record and summarize how bioinformatics researchers analyze tumor tissue slices, ensuring their visualization aligns with their real-world tasks. Experts might use LLMs to update their knowledge of a certain domain or generate drafts for academic papers based on high-level summaries. Additionally, LLMs support communication by helping researchers articulate their needs to domain experts and translate domain-specific terminology into plain language.
@@
Simulator: For the simulator role, LLMs are often used to simulate responses to questionnaires, predict potential user feedback, or test outcomes of workflows. By simulating user behavior, LLMs help researchers uncover hidden use cases, evaluate system usability, and even identify cognitive load issues before presenting the system to real users. Additionally, LLMs can take on the role of a critical reviewer, challenging the reasoning and logic behind a system and suggesting potential improvements. For instance, a proficient PhD student might use an LLM to refine their system by simulating user responses during the evaluation process before they actually present the system to target users for testing. This role is most frequently relied upon by proficient PhD students who are experienced enough to build simulated insights into their design studies effectively.
@@
Programmer: The programmer role is a commonly used applications of allowing LLMs to help generate, optimize, and revise code, streamlining the technical implementation process from prototyping to debugging. Specifically, researchers at varying levels of expertise, including novices, intermediates, and proficient researchers, reported using LLMs for programming support. For novices, LLMs provide guidance for setting up their code frameworks or generating simple scripts, such as creating mock datasets or pre-processing data. Intermediate researchers use LLMs for more complex coding tasks, such as creating JavaScript code by translating hand-drawn drafts into a simple visualization, while proficient researchers are able to leverage LLMs to simplify SVG elements or optimize a visualization’s runtime performance.
@@
Assistant: The assistant role refers to the general-purpose support LLMs offer to boost researchers’ productivity by automating repetitive or time-consuming tasks. Common users include searching for academic papers, summarizing a collaborator’s background from a CV, or generating scripts from user study recordings. For example, a novice researcher might use an LLM to summarize a complex academic paper into key points they can easily understand. A postdoc researcher could use it to transcribe interview recordings and distill key takeaways from them. Unlike the other roles, this role is used by researchers across all experience levels.